U Elock cannot run at such higha ih the same as it did at sea level. the speed of the water in the 1-em pipe, the dlameter pirsped n pipe the se comnected to a 2-em diameter pipe. Compared to e speed in the 2-cm pipe is: a) one-quarter b) one-hal c) the same d) double e) four times 13. An nl
Answer #1

12. Q = A1 v1 = A2 v2

and A = pi D^2 / 4

(pi D1^2 / 4) v1 = (pi D2^2 / 4) v2

v1 / v2 = (D2/D1)^2 = (2/1)^2 = 4

so v2 = v1/ 4

Ans: (a) one-quarter